In 5 years, Alessandro will eb three times as old as Frederico will be then. The sum of their ages now is 86. How old will each of them be in 3 years?
Answer by chana(69)

In 5 years, Alessandro will eb three times as old as Frederico will be then. The sum of their ages now is 86. How old will each of them be in 3 years?
sol:
let ales age = x
fred age = y
in 5 years,
(x+5) = 3(y+5)---------- 1
present;
x + y = 86 ---------- 2
solve 1 & 2 as equations which gives x = 67 , y = 19
so after 3 years , x = 70 , y = 22